### v2

@@ C.2   Repeated cross sections
We also present the influence function of our STS estimator when repeated cross section data are available. [^9] The derivation here is essentially the same in de Chaisemartin and D’Haultfœuille (2018), who present the influence function of the Wald-DID estimator in repeated cross section settings with two periods and two groups.. Let $\theta_{e,l}^{r}$ define
@@
In repeated cross section settings, our STS estimator is the sample analog of $\theta_{e,l}^{r}$, which we denote $\hat{\theta}_{e,l}^{r}$. By the similar arguments in the proof of Theorem 4, one can show that $\hat{\theta}_{e,l}^{r}$ is consistent for $CLATT_{e,e+l}$ under Assumptions 7, 10, 11, 14, 19, and 20, and have the following influence function:
@@ C.3   Triple DID-IV designs
In panel data settings, our triple DID-IV estimator is the sample analog of $\theta_{e,l,g}^{p}$ and has the following influence function:
@@
In repeated cross section data settings, our triple DID-IV estimator is the sample analog of $\theta_{e,l,g}^{r}$ and has the following influence function:
@@ Appendix D Decomposing the TWFEIV estimand in Oreopoulos (2006)
In this appendix we quantify the bias terms in the TWFEIV estimand, arising from the bad comparisons in Oreopoulos (2006). Hereafter, let $e=1947$ and $U=1957$ denote England and North Ireland respectively, and let $N_{e,t}$ denotes the sample size for England in cohort $t$. Let $R_{e,t}$ denote the sample mean of random variable $R_{i,t}$ in region $e$ ($e=1947,1957$) and cohort $t$:
@@
By the similar arguments in the proof of Theorem 4, one can show that in staggered DID-IV designs, we have
@@
Combining the result (30) with (34) and by the Slutsky’s theorem, we obtain the following decomposition result for the TWFEIV estimand in Oreopoulos (2006):
@@
where $n_{e,t}$ and $E[\hat{Z}_{i,t}|E_{i}=e]$ represent the population share and population residuals for England in cohort $t$ respectively.
@@
Figure 4: Decomposition result for the TWFEIV estimand in Oreopoulos (2006). Notes: The figure plots the estimated weights and the corresponding estimates for each $CLATT_{e,t}$ and $\Delta_{e,t}$ (bias term). The closed squares are the weights and the corresponding estimates for each $CLATT_{e,t}$. The closed triangles are the weights and the corresponding estimates for each $\Delta_{e,t}$ (bias term).
@@ Appendix E Comparing DID-IV to Fuzzy DID
E.1   Comparing dCDH’s identifying assumptions to ours
In this section, we investigate the detailed connections between dCDH’s identifying assumptions and ours. Because dCDH consider the repeated cross section settings, when we compare the identifying assumptions between the two papers, we will refer to Assumptions 14-18 in this article. For now, we ignore the difference in the definition of target parameter between dCDH and this article.
dCDH consider the following identification assumptions for the Wald-DID estimand to capture their target parameter, the SLATET [^12] In repeated cross section settings, the formal definition is provided in Assumption 29.. We note that in dCDH, a group variable $G$ plays the role of instrument $Z$ in our notation [^13] This follows from two observations in dCDH. First, dCDH call the groups $G=1$ and $G=0$ treatment and control groups, respectively. Second, the group variable $G$ is included in Assumption 26 (Treatment participation equation).. In the following, we therefore replace $G$ with $Z$ and call the groups $Z=1$ and $Z=0$ exposed and unexposed groups, respectively. [^14] Note that in Appendix B.2 in this article, we construct the group variable $E$ from the instrument path $Z$ in order to uncover Assumption 16 (No anticipation in the first stage). For the detail explanations for each assumption, see dCDH.
@@ E.1   Comparing dCDH’s identifying assumptions to ours
In Assumption 28, we modify the dCDH’s notation in order to ease the interpretation: we define $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$ to be the treatment status under time $T=t$ in group $Z$, which dCDH define as $D(t)$.
We now state the relationships between dCDH’s assumptions and ours. We first explain the similarities. First, both dCDH and this article impose the condition (39) in Assumption 24 (Fuzzy design), as Assumptions 15-17 in this article imply this condition [^15] From Assumptions 15-17, we have: $E[D|E=1,T=1]-[D|E=1,T=0]-(E[D|E=0,T=1]-E[D|E=0,T=0])$ $=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1,T=1]$ $=$ $Pr(CM^{Z}|E=1,T=1)>0.$
Next, the treatment participation equation in Assumption 26 corresponds to Assumption 15 (Monotonicity) in this article. The equivalence between the treatment participation equation and the potential treatment choices framework with Assumption 15 (Monotonicity) follows from the result of Vytlacil (2002). Here, we point out that dCDH implicitly assume the sharp assignment of the instrument $Z$ in their proofs: they assume $v_{10}=v_{00}$ in Assumption 26, that is, they assume that the thresholds in the treatment participation equation are the same between exposed and unexposed groups in period 00.
Finally, dCDH implicitly impose Assumption 14 (Exclusion Restriction) and Assumption 16 (No anticipation in the first stage) in this article when they formalize potential outcomes and treatment participation equations. Specifically, dCDH do not introduce the instrument path in their framework and define potential outcomes as $Y(D)$, which implies Assumption 14. As one can see in Assumption 26, dCDH assume that the treatment participation equation only depends on the current instrument status, which implies Assumption 16 since dCDH assume the sharp assignment of the instrument: $v_{10}=v_{00}$.
We next describe the differences between dCDH’s assumptions and ours. First, dCDH assume the parallel trends assumption in the untreated outcome (Assumption 27), whereas we assume the parallel trends assumption in the unexposed outcome. Second, dCDH assume the stable treatment rate assumption in an unexposed group (Assumption 25), whereas we assume the parallel trends assumption in the treatment between exposed and unexposed groups. Third, dCDH impose the independence assumption between unobserved heterogeneity $V$ and time variable $T$ given $Z$ in Assumption 26, whereas we do not impose such restriction. Forth, dCDH argue that the Wald-DID estimand requires the stable treatment effect assumption during the two periods (Assumption 28) to capture the SLATET, whereas we do not require such restriction in order for this estimand to identify the LATET. Finally, dCDH assume the condition (38) in Assumption 24 (Fuzzy design), which requires that the treatment rate in an exposed group increases between the two periods, whereas we do not assume this condition.
@@
When Assumption 25 is violated, dCDH consider the following assumption in order for the Wald-DID estimand to capture the SLATET.
@@
where we define161616dCDH denote SLATET and SLATET’ as $\Delta$ and $\Delta^{\prime}$ respectively. When we define the SLATET, we use $D_{0}(0)=D_{0}(1)$, which follows from the sharp assignment of the instrument: $v_{10}=v_{00}$.:
@@ E.2   Strong and asymmetric restrictions in Fuzzy DID designs
In this section, we show that dCDH’s identifying assumptions implicitly impose strong and asymmetric restrictions on treatment adoption behavior across units between exposed and unexposed groups. To do so, we first prove some lemmas implied by dCDH’s assumptions. Based on these lemmas, we then create the tables as in Table 1 and Table 2 in section 2.4 in this article, which present all the types in exposed and unexposed groups respectively in Fuzzy DID designs, and confirm that dCDH’s assumptions ex ante exclude many types in exposed and unexposed groups and its restrictions is heavily assigned on the latter group. At the end of this section, we provide empirical implications of the distinction between dCDH’s restrictions and ours in this article.
@@
Lemma 1 shows that in Fuzzy DID designs, the treatment status changes between period 00 and period $1$ in a monotone (uniform) way in exposed and unexposed groups, respectively. We note that the treatment participation equation itself does not imply Assumption 30: the independence between the unobserved heterogeneity $V$ and time variable $T$ given $Z$ yields this assumption. Indeed, in this article, we do not assume this restriction.
@@
Assumption 31 (Monotonicity Assumption w.r.t. time in an unexposed group).
@@
Assumption 31 requires that time affects the treatment adoption process from period 00 to period $1$ in a monotone (uniform) way in an unexposed group $z=0$. In terms of the notation introduced in section 2.4 in this article, this assumption implies that the group variable $G^{T}$ can take three values with non-zero probability. Hereafter, we consider the type of Assumption 31 which rules out the existence of the time-defiers $DF^{T}$ in an unexposed group.
@@
The first equivalence follows from $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 26. Equation (40) implies that there are no time comliers $CM^{T}$ in an unexposed group ($z=0$). ∎
Finally, we show that the condition (38) in Assumption 24 (Fuzzy design) and Assumption 26 imply that in an exposed group ($z=1$), there are no units switching from treatment to non treatment between period 00 and period $1$. In other words, in Assumption 30, we should have:
@@
We now clarify how dCDH’s identifying assumptions restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively. Table 9 and Table 10 show that their identifying assumptions exclude many types in both groups and the restrictions are heavily imposed on an unexposed group [^17] In both tables, we can only observe either $D_{0}(z)$ or $D_{1}(z)$ in group $z$ as we consider repeated cross section settings.: there exist five and three types in exposed and unexposed groups respectively, where the types painted in gray color are excluded in both groups.
The distinction between dCDH’s restrictions and ours is apparent when we compare Table 9 and Table 10 to Table 1 and Table 2 in this article [^18] We note that Tables 1-2 in this article is constructed under panel data settings, but the types represented in these tables remain consistent even when considering repeated cross section settings. The only modification is that as in Tables 9-10, one can only observe either $D_{0}(z)$ or $D_{1}(z)$ for unit $i$ in group $z$ in each table.: in both groups, we only exclude the defiers $DF^{Z}$ by Assumption 15 (Monotonicity), whereas dCDH additionally exclude the type $NT^{Z}\land DF^{T}$ in an exposed group by Lemma 3, and additionally exclude the time defiers $DF^{T}$ and the time compliers $CM^{T}$ in an unexposed group by Lemma 2.
@@
These differences between dCDH’s restrictions and ours seem critical in many empirical settings: unlike the monotonicity assumption, which eliminate the compliers $CM^{Z}$ in both groups, dCDH’s additional restrictions seem difficult to assess from the institutional knowledge in a given application, thus the practitioners may want to avoid imposing these restrictions in empirical work. For instance, in some cases, we may have no valid reason for excluding the specific type $NT^{Z}\land DF^{T}$ in an exposed group, who are not affected by instrument but affected by time. Similarly, in other cases, we may have no justification to exclude the time defiers $DF^{T}$ and the time compliers $CM^{T}$ in an unexposed group because of the steady growth of the treatment rate.
Although these restrictions in dCDH seem strong, one might argue that each assumption, such as the condition (38) in Assumption 24 and Assumption 25 (Stable percentage of treated units in an unexposed group), can be justifiable in certain applications. However, these assumptions give rise to two additional issues. First, when we impose the condition (38) in Assumption 24, we should test the plausibility of this condition from the data in practice because this condition excludes the case that the treatment rate in an exposed group is stable during the two periods: $E[D|Z=1,T=1]=E[D|Z=1,T=0]$. Although dCDH argue that ”Assumption $1$ (Assumption 24 in this article) is just a way to define the treatment and the control group (exposed and unexposed group in this article) in our fuzzy setting”, they do not explain how we should proceed when this condition is not met in practice.
Second, when we impose Assumption 25, we should find or estimate the groups whose treatment rates are stable over time, and this complicates our estimation procedure in practice. Although dCDH propose the two-step estimation method when we should estimate the stable treatment group from the data, its performance and feasibility is not clear, especially when we have multiple periods and multiple groups.
@@ E.3   Decomposing the SLATET
In the previous section, we investigate how dCDH’s identifying assumptions ex ante restrict the treatment adoption behavior across units between exposed and unexposed groups. In this section, under these restrictions, we show the decomposition result for dCDH’s target parameter, the SLATET, and state the concerning implications.
@@
Theorem 7 shows that we can decompose the SLATET into a weighted average of two different causal parameters under Fuzzy DID designs: one parameter measures the treatment effects among the time compliers $CM^{T}$ in an exposed group and the other parameter measures the treatment effects among the type $(CM^{Z}\land NT^{T})$ in an exposed group.
This theorem has two concerning implications. First, the interpretation of the SLATET is less clear than that of the LATET because it consists of two different causal parameters. Although the SLATET indeed captures the ”local” average treatment effect, it may be difficult to interpret economically in a given application because one of the two causal parameters captures the treatment effects, for those who are induced to treatment by time.
Second, the SLATET may not be policy-relevant even when the researchers exploit the policy change of their interest as an instrument for the treatment. Suppose that the instrument represents the important policy lever for the researchers, and they are intrinsically interested in the treatment effects among the compliers $CM^{Z}$ in an exposed group. In this situation, the LATET is a policy-relevant parameter by construction. On the other hand, the SLATET only captures the treatment effects among the type $(CM^{Z}\land NT^{T})$, which is the sub-population of the compliers $CM^{Z}$, even when the proportion of the time compliers $CM^{T}$ is zero within that group. As the fraction of the time compliers $CM^{T}$ increases in that group, the SLATET puts more weight on the treatment effects among the time compliers, and this would contaminate the interpretation of the magnitude of this parameter even in such policy-relevant context.
Using the decomposition result for the SLATET, we finally clarify the interpretation of Assumption 29. When Assumption 25 (Stable percentage of treated units in an unexposed group) is violated, dCDH additionally consider Assumption 29 in order for the Wald-DID estimand to identify the SLATET, and call this the ”homogeneous” treatment effect assumption between exposed and unexposed groups.
We show that this assumption can not be interpreted as requiring the ”homogeneous” treatment effect among the same sub-population between exposed and unexposed groups. To do so, we first note that if we do not assume Assumption 25, we have Table 11 in an unexposed group instead of Table 10 in Appendix E.2. Here, we assume that Assumption 26 and 31 exclude the defiers $DF^{Z}$ and the time defiers $DF^{T}$ in an unexposed group, respectively.
@@
By combining equations (42) and (43), we can conclude that Assumption 29 ($STATET=SLATET^{\prime}$) does not require the homogeneous treatment effects in the same sub-population between the two groups. Rather, this assumption requires that a weighted average of the treatment effects among the time compliers $CM^{T}$ and the type $(CM^{Z}\land NT^{T})$ in an exposed group should be equal to the treatment effect among the time compliers $CM^{T}$ in an unexposed group. This interpretation implies that the requirement of the ”homogeneous” treatment effect in Assumption 29 is less clear than one might expect.

### v4

@@ C.2   Repeated cross sections
We also present the influence function of our STS estimator when repeated cross section data are available. [^10] The derivation here is essentially the same in de Chaisemartin and D’Haultfœuille (2018), who presented the influence function of the Wald-DID estimator in repeated cross section settings with two periods and two groups.. Let $\theta_{e,l}^{r}$ define
@@
In repeated cross section settings, our STS estimator is the sample analog of $\theta_{e,l}^{r}$, which we denote $\hat{\theta}_{e,l}^{r}$. Using similar arguments in the proof of Theorem 4, one can show that $\hat{\theta}_{e,l}^{r}$ is consistent for $CLATT_{e,e+l}$ under Assumptions 7, 10, 11, 14, 19, and 20, and have the following influence function:
@@ C.3   Triple DID-IV designs
In panel data settings, our triple DID-IV estimator is the sample analog of $\theta_{e,l,g}^{p}$ and has the following influence function.
@@
In repeated cross section data settings, our triple DID-IV estimator is the sample analog of $\theta_{e,l,g}^{r}$ and has the following influence function.
@@ Appendix D Decomposing the TWFEIV estimand in Oreopoulos (2006)
Here, we quantify the bias terms in the TWFEIV estimand, arising from the bad comparisons in Oreopoulos (2006). Let $e=1947$ and $U=1957$ denote England and North Ireland respectively, and let $N_{e,t}$ denotes the sample size for England in cohort $t$. Let $R_{e,t}$ denote the sample mean of random variable $R_{i,t}$ in region $e$ ($e=1947,1957$) and cohort $t$:
@@
Using similar arguments in the proof of Theorem 4, one can show that in staggered DID-IV designs, we have
@@
Combining the result (30) with (34) and by Slutsky’s theorem, we obtain the following decomposition result for the TWFEIV estimand in Oreopoulos (2006):
@@
where $n_{e,t}$ and $E[\hat{Z}_{i,t}|E_{i}=e]$ represent the population share and residuals for England in cohort $t$ respectively.
@@
Figure 4: Decomposition result for the TWFEIV estimand in Oreopoulos (2006). Notes: This figure plots the estimated weights and the corresponding estimates for each $CLATT_{e,t}$ and $\Delta_{e,t}$ (bias term). The closed squares are the weights and the corresponding estimates for each $CLATT_{e,t}$. The closed triangles are the weights and the corresponding estimates for each $\Delta_{e,t}$ (bias term).
@@ Appendix E Comparing DID-IV to Fuzzy DID
E.1   Comparing the identifying assumptions in dCDH and this article
In this section, we investigate the detailed connections between dCDH’s identifying assumptions and ours. Because dCDH consider the repeated cross section settings, when we compare the identifying assumptions between the two papers, we refer to Assumptions 14-18. For now, we ignore the difference in the definition of target parameter between dCDH and this article.
dCDH consider the following identification assumptions for the Wald-DID estimand to capture their target parameter, the SLATET [^13] In repeated cross section settings, the formal definition is provided in Assumption 29.. In dCDH, a group variable $G$ plays the role of instrument $Z$ in our notation [^14] This follows from two observations in dCDH. First, dCDH call the groups $G=1$ and $G=0$ treatment and control groups, respectively. Second, the group variable $G$ is included in Assumption 26 (Treatment participation equation).. We therefore replace $G$ with $Z$ and call the groups $Z=1$ and $Z=0$ exposed and unexposed groups, respectively. [^15] Note that in Appendix B.2, we construct the group variable $E$ from the instrument path $Z$ to uncover Assumption 16 (No anticipation in the first stage), the hidden assumption in the previous literature. For the detailed explanations for each assumption, see dCDH.
@@ E.1   Comparing the identifying assumptions in dCDH and this article
In Assumption 28, we modify the dCDH’s notation $D(t)$ to ease its interpretation: we define $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$ to be the treatment status under time $T=t$ in group $Z$ instead of defining it as $D(t)$.
We now state the relationships between dCDH’s assumptions and ours. We first explain the similarities. First, both dCDH and this article impose the condition (39) in Assumption 24 (Fuzzy design), as Assumptions 15-17 imply this condition [^16] From Assumptions 15-17, we have: $E[D|E=1,T=1]-[D|E=1,T=0]-(E[D|E=0,T=1]-E[D|E=0,T=0])$ $=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1,T=1]$ $=$ $Pr(CM^{Z}|E=1,T=1)>0.$
Next, the treatment participation equation in Assumption 26 corresponds to Assumption 15 (Monotonicity) in this article. The equivalence between the treatment participation equation and the potential treatment choices framework with Assumption 15 (Monotonicity) follows from the result of Vytlacil (2002). We point out that dCDH implicitly assume the sharp assignment of the instrument $Z$ in their proofs: they assume $v_{10}=v_{00}$ in Assumption 26; that is, the thresholds in the treatment participation equation are the same between exposed and unexposed groups in period 00.
Finally, dCDH implicitly impose Assumption 14 (Exclusion restriction) and Assumption 16 (No anticipation in the first stage) when they formalize potential outcomes and treatment participation equations. Specifically, dCDH do not introduce the instrument path in their framework and define potential outcomes as $Y(D)$, which implies Assumption 14. As one can see in Assumption 26, dCDH assume that the treatment participation equation only depends on the current instrument status, which implies Assumption 16 since dCDH assume the sharp assignment of the instrument: $v_{10}=v_{00}$.
The differences between dCDH’s assumptions and ours are as follows. First, dCDH assume the parallel trends assumption in the untreated outcome (Assumption 27), whereas we assume the parallel trends assumption in the unexposed outcome. Second, dCDH assume the stable treatment rate assumption in an unexposed group (Assumption 25), whereas we assume the parallel trends assumption in the treatment between exposed and unexposed groups. Third, dCDH impose the independence assumption between unobserved heterogeneity $V$ and time variable $T$ given $Z$ in Assumption 26, whereas we do not impose such restriction. Forth, dCDH argue that the Wald-DID estimand requires the stable treatment effect assumption during the two periods (Assumption 28) to capture the SLATET, whereas we do not require such restriction for this estimand to identify the LATET. Finally, dCDH assume the condition (38) in Assumption 24 (Fuzzy design), which requires that the treatment rate in an exposed group increases between the two periods, whereas we do not assume this condition.
@@
When Assumption 25 is violated, dCDH consider the following assumption for the Wald-DID estimand to capture the SLATET.
@@
where we define171717dCDH denote SLATET and SLATET’ as $\Delta$ and $\Delta^{\prime}$ respectively. When we define the SLATET, we use $D_{0}(0)=D_{0}(1)$, which follows from the sharp assignment of the instrument: $v_{10}=v_{00}$.:
@@ E.2   Strong and asymmetric restrictions in Fuzzy DID designs
We show that dCDH’s identifying assumptions implicitly impose strong and asymmetric restrictions on treatment adoption behavior across units between exposed and unexposed groups. First, we prove some lemmas implied by dCDH’s assumptions. Based on these lemmas, we then create the tables that present all the types in exposed and unexposed groups under Fuzzy DID designs. These tables show that dCDH’s assumptions ex ante exclude many types in both groups and these restrictions are heavily assigned on an unexposed group. At the end, we provide the empirical implication of the distinction between dCDH’s restrictions and ours in DID-IV designs.
@@
Lemma 1 shows that in Fuzzy DID designs, the treatment status changes between period 00 and period $1$ in a monotone (uniform) way in exposed and unexposed groups, respectively. Note that the treatment participation equation itself does not imply Assumption 30: the independence between the unobserved heterogeneity $V$ and time variable $T$ given $Z$ yields this assumption. Indeed, in this article, we do not assume this restriction.
@@
Assumption 31 (Monotonicity assumption with respect to time in an unexposed group).
@@
Assumption 31 requires that time affects the treatment adoption process from period 00 to period $1$ in a monotone (uniform) way in an unexposed group $z=0$. In terms of the notation introduced in Subsection 2.4, this assumption implies that the group variable $G^{T}$ can take three values with non-zero probability. Hereafter, we consider the type of Assumption 31 which rules out the existence of the time-defiers $DF^{T}$ in an unexposed group.
@@
The first equivalence follows from $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 26. Equation (40) implies that there are no time compliers $CM^{T}$ in an unexposed group ($z=0$). ∎
Finally, we show that the condition (38) in Assumption 24 (Fuzzy design) and Assumption 26 imply that in an exposed group ($z=1$), there are no units switching from treatment to non-treatment between period 00 and period $1$. In other words, in Assumption 30, we should have:
@@
We now clarify how dCDH’s identifying assumptions restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively. Tables 9 - 10 show that their identifying assumptions exclude many types in both groups and the restrictions are heavily imposed on an unexposed group [^18] In both tables, we can only observe either $D_{0}(z)$ or $D_{1}(z)$ in group $z$ as we consider repeated cross section settings.: five and three types exist in exposed and unexposed groups respectively, where the types painted in gray color are excluded in both groups.
The distinction between dCDH’s restrictions and ours is apparent when we compare Tables 9 - 10 to Table 1 - 2191919Tables 1-2 are constructed under panel data settings, but the types represented in these tables remain consistent even when considering repeated cross section settings. The only modification is that as in Tables 9-10, one can only observe either $D_{0}(z)$ or $D_{1}(z)$ for unit $i$ in group $z$ in each table.: in both groups, we only exclude the defiers $DF^{Z}$ by Assumption 15 (Monotonicity), whereas dCDH additionally exclude the type $NT^{Z}\land DF^{T}$ in an exposed group by Lemma 3, and additionally exclude the time defiers $DF^{T}$ and the time compliers $CM^{T}$ in an unexposed group by Lemma 2.
@@
These differences between dCDH’s restrictions and ours seem critical in many empirical settings. Unlike the monotonicity assumption, which eliminate the compliers $CM^{Z}$ in both groups, dCDH’s additional restrictions seem difficult to assess from the institutional knowledge in a given application, and the practitioners may want to avoid imposing these restrictions in empirical work. For instance, in some cases, we may have no valid reason for excluding the specific type $NT^{Z}\land DF^{T}$ in an exposed group, who are not affected by instrument but affected by time. Similarly, in other cases, we may have no justification to exclude the time defiers $DF^{T}$ and time compliers $CM^{T}$ in an unexposed group because of the steady growth of the treatment rate.
Although these restrictions in dCDH seem strong, one might argue that each assumption, such as the condition (38) in Assumption 24 and Assumption 25 (Stable percentage of treated units in an unexposed group), can be justifiable in certain applications. However, these assumptions give rise to two additional issues. First, when we impose the condition (38) in Assumption 24, we should test its plausibility from the data because this condition excludes the case that the treatment rate in an exposed group is stable during the two periods: $E[D|Z=1,T=1]=E[D|Z=1,T=0]$. Although dCDH argue that ”Assumption $1$ (Assumption 24 in this article) is just a way to define the treatment and the control group (exposed and unexposed group) in our fuzzy setting”, they do not explain how we should proceed when this condition is violated in practice.
Second, when we impose Assumption 25, we should find or estimate the groups whose treatment rates are stable over time. Although dCDH propose the two-step estimation method when we should estimate the stable treatment group from the data, its performance and feasibility are not clear, especially for multiple period and multiple group settings.
@@ E.3   Decomposing the SLATET
In the previous subsection, we investigate how dCDH’s identifying assumptions ex ante restrict the treatment adoption behavior across units between exposed and unexposed groups. In this subsection, under these restrictions, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters.
@@
Theorem 7 shows that we can decompose the SLATET into a weighted average of two different causal parameters under Fuzzy DID designs: one parameter measures the treatment effects among the time compliers $CM^{T}$ in an exposed group and the other measures the treatment effects among the type $(CM^{Z}\land NT^{T})$ in an exposed group.
This theorem has two concerning implications. First, the interpretation of the SLATET is less clear than that of the LATET because it consists of two different causal parameters. Although the SLATET captures the ”local” average treatment effect, it may be difficult to interpret economically in a given application because one of the two causal parameters captures the treatment effects, for those who are induced to treatment by time.
Second, the SLATET may not be policy-relevant even when the researchers exploit the policy change of their interest as an instrument for treatment. Suppose that the instrument represents an important policy lever for researchers, who are intrinsically interested in the treatment effects among the compliers $CM^{Z}$ in an exposed group. In this situation, our target parameter, the LATET, is a policy-relevant parameter by construction. By contrast, the SLATET only captures the treatment effects among the type $(CM^{Z}\land NT^{T})$, which is the sub-population of the compliers $CM^{Z}$, even when the proportion of the time compliers $CM^{T}$ is zero within that group. As the share of the time compliers $CM^{T}$ increases, the SLATET puts more weight on the treatment effects among the time compliers, which contaminates the interpretation of the magnitude of this parameter even in such policy-relevant contexts.
Using the decomposition result for the SLATET, we finally clarify the interpretation of Assumption 29. When Assumption 25 (Stable percentage of treated units in an unexposed group) is violated, dCDH additionally consider Assumption 29 for the Wald-DID estimand to identify the SLATET, and call this the ”homogeneous” treatment effect assumption between exposed and unexposed groups.
However, we show that this assumption cannot be interpreted as requiring the ”homogeneous” treatment effect among the same sub-population between exposed and unexposed groups. If we do not assume Assumption 25, we have Table 11 in an unexposed group instead of Table 10 in Appendix E.2. Here, we assume that Assumption 26 and 31 exclude the defiers $DF^{Z}$ and the time defiers $DF^{T}$ in an unexposed group, respectively.
@@
By combining equations (42) and (43), we can conclude that Assumption 29 ($STATET=SLATET^{\prime}$) does not require the homogeneous treatment effects in the same sub-population between the two groups. Rather, this assumption requires that a weighted average of the treatment effects among the time compliers $CM^{T}$ and the type $(CM^{Z}\land NT^{T})$ in an exposed group should be equal to the treatment effect among the time compliers $CM^{T}$ in an unexposed group. This interpretation implies that the requirement of the ”homogeneous” treatment effect in Assumption 29 is less clear than what one might expect.